# Perception-aware Sampling for Scatterplot Visualizations

### v1

@@ 6.1. Perceptual quality of sampling methods
User study. To evaluate PAwS, we conducted a user study assessing its efficacy across three key aspects: (1) preserving density variations, (2) retaining outliers and overall data shape, and (3) maintaining correlation trends, clusters, and other patterns.
[…]
Figure 15 presents aggregated results across all datasets and participants: i.e., $\sim 50$ trials per sample pair. For each sample pair—comparing a PAwS sample to an alternative sample—we report how often participants preferred each method across all trials. PAwS shows strong performance across all four evaluation criteria, while Blue-noise is the method ranked closer to it in terms of user preference. However, Blue-noise is computationally expensive and often fails to preserve data shape—most notably on Hidden Correlation—leading participants to favor PAwS there. Probabilistic-based methods generally rank below PAwS, but we observed participants only decisively prefer PAwS when those methods fail to capture shape and outliers (as is the case on Hidden Correlation and ACSI). Finally, while PAwS outperforms diversity‑based methods, Max-Min produces nearly identical samples in some datasets, frequently making it hard for participants to choose between them. We show a per-dataset analysis of the results in Appendix A.
Runtime performance. Finally, we discuss the runtime performance of the algorithms on the Hidden Correlation dataset ($\sim 3.5M$), which is the largest one in our evaluation (Figure 14, right). We note that PAwS and Max-Min are the fastest among algorithms with reasonable perceptual efficacy. Their runtimes are similar, because they implement the same algorithm with modified objectives. While Random and DBS are faster, they typically produce samples of poor perceptual quality. The KNN algorithm, which is used by DBS to identify points in sparse areas, dominates its running time, which stays constant across sample sizes. The runtimes of Vas and Blue-noise are significantly longer, by 2–3 orders of magnitude. We note that Vas takes a long time to converge (${\sim}23$ minutes for a sample of 250 points), but we have observed that it produces samples of similar quality after one pass over the data (${\sim}5$ minutes for a sample of 250 points). Figure  14 reports both runtimes (till full convergence and for a single pass). We note that Blue-noise is extremely inefficient (${\sim}3.6$ hours for a sample of 250 points), which aligns with prior analyses (Quadri et al., 2022; Chen et al., 2019).
Key takeaways: PAwS robustly outperforms state-of-the-art sampling methods, with emphasized gains in smaller sample sizes. It is also orders of magnitude faster than the other state-of-the-art. Probabilistic methods have consistently low performance. While Blue-noise often achieves good perceptual similarity, its runtime is prohibitive in large datasets. Diversity-based methods have more stable behavior and achieve similar performance to PAwS for large sample sizes, but Vas is again significantly slower. A user study further confirms that PAwS samples are frequently preferred by humans.
6.2. Performance of approximate visualization
In this section, we evaluate the effectiveness of approximate visualizations generated by ApproPAwS (Algorithm 2), using our perception-aware compression (Section 4.1). Compression splits the canvas area into non-overlapping bounding boxes based on two thresholds, $\lambda$ and $\sigma$, controlling data approximation and homogeneity of perception weights, respectively. These parameters affect the level of compression: lower thresholds result in more and smaller boxes (lower compression), thus, sample points drawn randomly within a box match better the underlying data and lead to lower distortion (example in Figure 10(a)). We present results at three compression levels: (1) low: $\langle\lambda,\sigma\rangle=\langle 0.001,0.001\rangle$, (2) medium: $\langle\lambda,\sigma\rangle=\langle 0.002,0.001\rangle$, and (3) high: $\langle\lambda,\sigma\rangle=\langle 0.003,0.01\rangle$. For consistency and ease of exposition, we use the same threshold values for all datasets but ACSI. For ACSI, to observe notable differences across compression schemes, we vary the thresholds as: (1) low: $\langle\lambda,\sigma\rangle=\langle 0.001,0.001\rangle$, (2) medium: $\langle\lambda,\sigma\rangle=\langle 0.005,0.001\rangle$, and (3) high: $\langle\lambda,\sigma\rangle=\langle 0.01,0.01\rangle$. We emphasize that the derived representations may differ in the number of boxes: e.g., hidden correlation has 9,658, 4,042, and 993 boxes across the three levels (Figure 10(a)), while estate correlation has 3,491, 1,184, and 448 boxes. We include additional details in Appendix A. In practice, the thresholds can be tuned by data analysts using a grid search approach to reach the desired level of compression.
@@ 8. Summary and Future Directions
In this paper, we introduce the problem of perception-aware sampling for scatterplot visualizations. We propose perception-augmented databases, which augment datasets with saliency information, and design novel sampling methodologies to derive perception-aware samples (PAwS) and approximate visualizations (ApproPAwS). Our evaluation demonstrated significant gains, both in the perceptual quality of the sample and sampling runtime.
Many interesting questions remain on targeting samples toward particular visual analysis tasks and optimizing samples to multiple target visualizations. On-the-fly sample augmentation can be useful to support interactivity in visualizations (e.g., when a user zooms in), and the biggest challenge is to achieve interactive sampling times. With the assistance of eye-tracking technologies, one may also aim for on-the-fly sample adaptation, based on user interactions and eye-gaze. Importantly, evaluating sampling methods in the context of visualizations is challenging, as there is no established benchmark. Creating such a benchmark would facilitate future research and make the comparison among different techniques easier and more robust.
@@ Appendix A Appendix
Overview of experimental design. Figure 18 provides an overview of the experimental design choices we described in Section 6 of the main paper. We show the aggregate saliency maps derived using our approach and describe in more detail our methodology for computing perceptual similarity. Figure 19 shows the perception weights for the datasets in our evaluation, assigning a higher perception value to highly salient or highly dense areas.
Perceptual similarity metrics. Figure 20 shows extensive results about the behavior of different sampling methods across six datasets and five evaluation metrics introduced in Section 5 of the main paper. The observations we make are consistent with the discussion in Section 6, and different metrics show similar results. PAwS has strong behavior across all datasets and metrics, while probabilistic-based approaches consistently underperform. Max-Min and Vas typically perform better as the sample size increases. The behavior of Blue-noise varies across datasets, while in hidden correlation performs as poorly as probabilistic-based methods with extremely low scores in SIM, CC, and 1-JSD metrics. We note EMD is a difficult metric to optimize for, and in some datasets (i.e., in epileptic seizure) all sampling methods reach lower scores.
@@
Qualitative samples and user study results per dataset. Figure 21 shows the visualizations we used in our user study assessing the efficacy of different sampling methods across three key aspects: (1) preserving density variations, (2) retaining outliers and overall data shape, and (3) maintaining correlation trends, clusters, and other patterns. Figures 22, 23, and 24 show the results per dataset and sampling pair —comparing a PAwS sample to an alternative sample—across these three key criteria. Figure 25 shows the result for the best overall sample per dataset. We observe PAwS is frequently preferred by humans compared to state-of-the-arts methods, while the preference becomes more evident in datasets where other sampling methods fail to represent some trend in the data; for instance, in hidden correlation, PAwS is overly favored compared to all other methods, which fail to represent both the correlation trend and the overall data distribution. Similarly, in ACSI, humans overly prefer PAwS compared to probabilistic-based approaches that create an evident “gap” in the sample.
@@
(a) Random
(b) Density-biased
(c) Vas
(d) Max-Min
(e) Blue-noise
@@
(a) Random
(b) Density-biased
(c) Vas
(d) Max-Min
(e) Blue-noise
@@
(a) Random
(b) Density-biased
(c) Vas
(d) Max-Min
(e) Blue-noise
@@
(a) Random
(b) Density-biased
(c) Vas
(d) Max-Min
(e) Blue-noise

### v3

@@ 8. Summary and Future Directions
In this paper, we introduce the problem of perception-aware sampling for scatterplot visualizations. We propose perception-augmented databases, which model perception as meta-data, and we design novel sampling methodologies to derive perception-aware samples (PAwS) and approximate visualizations (ApproPAwS). Our evaluation demonstrated significant quality and runtime gains, and a user study confirmed that humans prefer PAwS samples.
Many interesting questions remain on targeting samples toward particular visual analysis tasks and optimizing samples to multiple target visualizations. On-the-fly sample augmentation can be useful to support interactivity in visualizations (e.g., zooming in), and the biggest challenge is to achieve interactive sampling times. With eye-tracking technologies, one may also aim for on-the-fly sample adaptation, based on user interactions and eye-gaze. Importantly, evaluating sampling methods in the context of visualizations is challenging, as there is no established benchmark. Creating such a benchmark would facilitate future research and make the comparison among different techniques easier and more robust.
Acknowledgements.
This work was supported by Adobe seed funding, a Google DANI award, and by the NSF under grants IIS-2237585 and IIS-2311575.
@@ Appendix A Appendix
Overview of experimental design. Figure 18 provides an overview of the experimental design choices we described in Section 6. We show the aggregate saliency maps derived using our approach and describe our methodology for computing perceptual similarity. Figure 19 shows the perception weights for the datasets in our evaluation, assigning a higher perception value to highly salient or highly dense areas.
Perceptual similarity metrics. Figure 20 shows extensive results about the behavior of different sampling methods across six datasets and five evaluation metrics introduced in Section 5. The observations we make are consistent with the discussion in Section 6, and different metrics show similar results. PAwS has strong behavior across all datasets and metrics, while probabilistic-based approaches consistently underperform. Max-Min and Vas typically perform better as the sample size increases. The behavior of Blue-noise varies across datasets, while in hidden correlation performs as poorly as probabilistic-based methods with extremely low scores in SIM, CC, and 1-JSD metrics. We note EMD is a difficult metric to optimize for, and in some datasets (i.e., in epileptic seizure) all sampling methods reach lower scores.
@@
[…]
Figure 22 reports how often participants preferred PAwS over an alternative baseline concerning their efficacy in maintaining correlation trends, clusters, and other patterns; Figure 23 reports how often participants preferred PAwS over an alternative baseline concerning their efficacy in retaining outliers and overall data shape; and Figure 24 report users’ preference concerning density variations. Finally, Figure 25 shows the result for the best overall sample per dataset. We observe PAwS is frequently preferred by humans compared to state-of-the-arts methods, while the preference becomes more evident in datasets where other sampling methods fail to represent some trend in the data; for instance, in hidden correlation, PAwS is overly favored compared to all other methods, which fail to represent both the correlation trend and the overall data distribution. Similarly, in ACSI, humans overly prefer PAwS compared to probabilistic-based approaches that create an evident underrepresented region in the sample.
@@
(a) PAwS vs Random
(b) PAwS vs DBS
(c) PAwS vs Vas
(d) PAwS vs Max-Min
(e) PAwS vs Blue-noise
@@
(a) PAwS vs Random
(b) PAwS vs DBS
(c) PAwS vs Vas
(d) PAwS vs Max-Min
(e) PAwS vs Blue-noise
@@
(a) PAwS vs Random
(b) PAwS vs DBS
(c) PAwS vs Vas
(d) PAwS vs Max-Min
(e) PAwS vs Blue-noise
@@
(a) PAwS vs Random
(b) PAwS vs DBS
(c) PAwS vs Vas
(d) PAwS vs Max-Min
(e) PAwS vs Blue-noise